OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

First find the slope of the line that passes through the point (4,6) and (-1,-4) use this \[m = \frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

-7/-5 @mebs

OpenStudy (anonymous):

No, not quite... do this \[m = \frac{ 6 - (-4) }{ 4 - (-1) }\]

OpenStudy (anonymous):

10/5 @mebs

OpenStudy (anonymous):

2 @mebs

OpenStudy (anonymous):

yes so now we have a slope m = 2 we can use this \[y - 1 = 2(x +1)\] solve for y

OpenStudy (anonymous):

2x+3

OpenStudy (anonymous):

In general you have a point (-1,1) which is (xo , yo) \[y - yo = m (x - xo)\] yes you are correct @snewton29 y = 2x +3

OpenStudy (anonymous):

So the answer is 2x+3 for my whole answer @mebs

OpenStudy (anonymous):

No. the slope is 2. that is your whole answer.

OpenStudy (anonymous):

New question , What is the value of x in the solution to the following system of equations X-Y=-3 X+3Y=5

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

y = x +3 sub that into x +3y = 5 solve x + 3(x+3) = 5

OpenStudy (anonymous):

I have to go. @cambrige take over.

OpenStudy (anonymous):

How would u solve that @cambrige

OpenStudy (anonymous):

yup m here

OpenStudy (anonymous):

Just remember..slope of two parallel lines always equal... and slope of 2 perpendicular lines -ve multiplicative inverse!!

OpenStudy (anonymous):

so for example if slope of a line is m=3, slope of a line parallel to it m'=3 sope of a line perpendicular to it m"=-(1/3)

OpenStudy (anonymous):

In answer 2 ur second Q, U can use wat @mebs mentioned known as 1.substitution method or 2.elimination method

OpenStudy (anonymous):

In sub method use any one of the equations to obtain the value of one(say x) of the two variables in terms of the other. Substitute this value in the second linear eq. to get value of that variable(say x). the use any eq and substitute value of variable(here x) to get the variable of y!!

OpenStudy (anonymous):

Tell me wen u r online I'll xplain u the 2nd method!!Cheers!!